the smaller of two number is 3 less than the larger. Four times the smaller is equal to twice the larger. Find the numbers
Answer by mananth(16946)

let x be smaller
& y be larger
..
x= y-3
x-y=-3......................1
...
4x = 2y
/2
2x=y
...
plug the value of y in (1)
x-2x=-3
-x=-3
/-1
x=3
plug value of x in (1)
x-y =-3
3-y=-3
-y=-3+ -3
-y=-6
/-1
y=6
..
m.ananth@ hotmail.ca